OpenStudy (anonymous):

Use the quadratic formula to solve the equation. x^2 – 7x – 6 = 0 Show work please.

OpenStudy (anonymous):

x^2-7x-6=0?

OpenStudy (lilg132):

\[x = \frac{ -b \pm \sqrt{b ^{2} -4ac} }{ 2a}\]

OpenStudy (anonymous):

yes it is an exponent

OpenStudy (lilg132):

x2 = a -7x = b -6 = c x2 = 1 -7x = -7 -6 = -6

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]where a=1, b=-7, and c =-6 This is because your equation is in the standard form of ax^2 + bx + c = 0. So, now, just make the substitutions.

OpenStudy (anonymous):

Hmm, ok thanks, let me work this out and I will reply to what I got to check if it is right.

OpenStudy (anonymous):

Hint: after you make your substitutions, pay special attention to the expression within the radical (the square root sign). You will see that the numbet in the radical is positive, so you will have two real solutions. But your final solution will have a radical in it.

OpenStudy (anonymous):